### v2

Abstract.
In a recent paper (Math. Ann. 393 (2025), 1769–1795), Elorreaga et al. have obtained a complete characterization of the notion of a $h$-dichotomy for ordinary differential equations on a finite-dimensional space in terms of the notions of $h$-expansiveness and $h$-noncriticality. Their results extended the previous results of Coppel and Palmer, which dealt with exponential dichotomies. The main objective of this note is to extend the results of Elorreaga et al. to arbitrary invertible evolution families that act on Banach spaces. We emphasize that our approach is completely different and considerably simpler from the one developed by Elorreaga et al. It is based on the time-rescaling method introduced by Dragičević and Silva.
1. Introduction
The notion of an exponential dichotomy (essentially introduced by Perron [17]) plays a fundamental role in the qualitative theory of nonautonomous differential equations and dynamical systems. It can be described as a nonautonomous counterpart to the classical notion of hyperbolicity as it requires that the phase space of a linear dynamics splits (at every moment of time) into two directions: the stable and the unstable direction. Along the stable direction, dynamics exhibits exponential contraction forward in time, while along the unstable direction it exhibits the same property backward in time (which corresponds to the exponential expansiveness forward in time). We refer to the important monographs [1, 2, 3, 8, 18, 19] for a detailed exposition of various aspects of the theory of exponential dichotomies, including many applications.
Despite its importance, due to the flexibility of nonautonomous dynamics, it is fairly easy to construct broad classes of dynamics which exhibit behavior similar to exponential dichotomies but where the rates of contraction (resp. expansion) along stable (resp. unstable) directions are not necessarily exponential. To our knowledge, Martin Jr. [10], Muldowney [13] and Naulin and Pinto [14] were the first to systematically study such generalized dichotomies, in which the rates of contraction and expansion are prescribed by some general functions (growth rates).
In their recent paper [7] (for related earlier work see [20]), the authors have obtained important new characterizations of the notion of a (uniform) $h$-dichotomy for nonautonomous ordinary differential equations (which is a special case of a more general notion of the $(h,k)$-dichotomy introduced in [14]). These characterizations are given in terms of the newly introduced notions of uniform $h$-noncriticality and $h$-expansiveness, motivated by the notions of uniform noncriticality and exponential expansiveness studied by Coppel [2] and Palmer [15]. In particular, the results from [7] extend some of those obtained by Palmer [15] from exponential to general $h$-dichotomies. Their approach relied on the idea that to any $h$-dichotomy we can associate a totally ordered topological group. This idea first appeared in [16].
The main objective of the present paper is to obtain versions of the results of [7] for general invertible evolution families that act on Banach spaces. Although it is likely that one can adapt the approach from [7] to our setting, we prefer to use a different approach, which relies on the so-called time-rescaling. The basic idea is that the existence of an $h$-dichotomy of some evolution family $T(t,s)$ is equivalent to the existence of an exponential dichotomy of a new evolution family $T_{h}(t,s)$. This technique was developed in [5] (building on the earlier work [6] for polynomial dichotomies) for the case of discrete time, while in the present paper we developed it for continuous time. We stress that some aspects of this approach were hidden in the arguments of previous works (see, for example, [4]) but without a systematic treatment.
The use of the approach described in the previous paragraph enables us to deduce facts about $h$-dichotomies directly from those about exponential dichotomies.
The paper is organized as follows. In Section 2, we recall the basic notions that will appear throughout the paper. In Section 3, we explore the connection between exponential and $h$-dichotomies. Afterwards, in Section 4 we introduce the concepts of uniform $h$-noncriticality and $h$-expansiveness for an arbitrary invertible evolution family $T(t,s)$ and note that these can also be characterized in terms of uniform noncriticality and exponential expansiveness of $T_{h}(t,s)$. Finally, in Section 5 we obtain the main results of this paper. Firstly, we obtain the version of [15, Theorem 1] for invertible evolution families on Banach spaces (see Theorem 1) providing characterizations of exponential dichotomies. Using the other results of our paper, we then obtain the version of Theorem 1 for $h$-dichotomies as a simple corollary (see Theorem 2).
2. Preliminaries
Throughout this note $X=(X,\|\cdot\|)$ will be a Banach space. By $\mathcal{B}(X)$ we denote the space of all bounded linear operators on $X$ equipped with the operator norm, which we also denote by $\|\cdot\|$.
We begin by recalling the notion of an evolution family.
Definition 1.
Let $a_{0}\in\mathbb{R}\cup\{-\infty\}$. We say that $\mathcal{T}=\{T(t,s):\ t\geq s>a_{0}\}\subset\mathcal{B}(X)$ is an evolution family if the following holds:
  • $T(t,t)=\text{\rm Id}$ for $t>a_{0}$, where Id denotes the identity operator on $X$;
  • for $t\geq s\geq r>a_{0}$,
    $T(t,s)T(s,r)=T(t,r);$
  • for $s>a_{0}$ and $v\in X$, $t\mapsto T(t,s)v$ is continuous on $s,\infty)$.
In addition, if $T(t,s)$ is an invertible operator for each $t\geq s>a_{0}$, we say that $\mathcal{T}$ is an invertible evolution family.
Remark 1.
Let $\mathcal{T}=\{T(t,s):\ t\geq s>a_{0}\}\subset\mathcal{B}(X)$ be an invertible evolution family. We can define $T(t,s)$ for $a_{0}<t<s$ by
$T(t,s):=T(s,t)^{-1}.$
Thus, in this case, we can view $\mathcal{T}$ as a family $\{T(t,s):\ t,s>a_{0}\}$.
Definition 2.
A growth rate is any bijective increasing map $h\colon(a_{0},\infty)\to(0,\infty)$, where $a_{0}\in\mathbb{R}\cup\{-\infty\}$.
Definition 3.
Let $h\colon(a_{0},\infty)\to(0,\infty)$ be a growth rate and $\mathcal{T}=\{T(t,s):\ t\geq s>a_{0}\}\subset\mathcal{B}(X)$ an evolution family. We say that $\mathcal{T}$ exhibits $h$-bounded growth on an interval $J\subset(a_{0},\infty)$ if there exist $K,\mu>0$ such that
$\|T(t,s)\|\leq K\left(\frac{h(t)}{h(s)}\right)^{\mu},\quad\text{for $t,s\in J$ with $t\geq s$.}$ (2.1)
Definition 4.
Let $h\colon(a_{0},\infty)\to(0,\infty)$ be a growth rate and $\mathcal{T}=\{T(t,s):\ t\geq s>a_{0}\}\subset\mathcal{B}(X)$ an invertible evolution family. We say that $\mathcal{T}$ exhibits $h$-bounded decay on an interval $J\subset(a_{0},\infty)$ if there exist $K,\mu>0$ such that
$\|T(t,s)\|\leq K\left(\frac{h(s)}{h(t)}\right)^{\mu},\quad\text{for $t,s\in J$ with $t\leq s$.}$ (2.2)
Remark 2.
In the particular case where $h(t)=e^{t}$, the notion of an $h$-bounded growth coincides with the classical notion of a bounded growth (see [12, p.334]). In addition, in this case $h$-bounded decay will be called bounded decay.
Remark 3.
  1. (1)
    Assume that an evolution family $\mathcal{T}=\{T(t,s):\ t\geq s>a_{0}\}\subset\mathcal{B}(X)$ is generated by a nonautonomous linear equation
    $x^{\prime}=A(t)x\quad t>a_{0},$ (2.3)
    where $A\colon(a_{0},\infty)\to\mathcal{B}(X)$ is a continuous map. In this case, (2.1) implies that
    $\|x(t)\|\leq K\left(\frac{h(t)}{h(s)}\right)^{\mu}\|x(s)\|,\quad\text{for $t,s\in J$ with $t\geq s$,}$ (2.4)
    where $t\mapsto x(t)$ is any solution of (2.3). This follows immediately, taking into account that $T(t,s)x(s)=x(t)$. Conversely, take any $s\in J$, $v\in X$ and let $x\colon(a_{0},\infty)\to X$ be the solution of (2.3) with $x(s)=v$. From (2.4) we obtain
    $\|T(t,s)v\|\leq K\left(\frac{h(t)}{h(s)}\right)^{\mu}\|v\|,\quad\text{for $t\in J$ with $t\geq s$.}$
    Since $v$ and $s$ were arbitrary, we conclude that (2.1) holds. The same discussion applies in relation to (2.2).
  2. (2)
    We note that an evolution family $\mathcal{T}=\{T(t,s):\ t\geq s>a_{0}\}\subset\mathcal{B}(X)$ exhibits bounded growth if and only if there exist $C,T>0$ such that
    $\|T(t,s)\|\leq C,\quad\text{for $t,s\in J$ with $t\in[s,s+T]$}.$ (2.5)
    Indeed, it follows from (2.1) that (2.5) holds with $T=1$ and $C=Ke^{\mu}$. Conversely, suppose that (2.5) holds with $C>1$ (we can increase $C$ to achieve this) and take arbitrary $t,s\in J$ with $t\geq s$. Let $n\in\mathbb{N}_{0}$ be such that ts=nT+rt-s=nT+r with $0\leq r<T$. Then
    $\begin{split}\|T(t,s)\|&=\|T(s+nT+r,s)\|\\ &=\|T(s+nT+r,s+nT)T(s+nT,s)\|\\ &\leq C\|T(s+nT,s)\|\\ &\leq C^{n+1}\\ &=e^{\frac{t-s+T-r}{T}\ln C}\\ &\leq Ce^{(t-s)\frac{\ln C}{T}},\end{split}$
    which implies that (2.1) holds with $K=C$ and $\mu=\frac{\ln C}{T}>0$.
We introduce the notion of a $h$-dichotomy for evolution families.
Definition 5.
Let $h\colon(a_{0},\infty)\to(0,\infty)$ be a growth rate and $\mathcal{T}=\{T(t,s):\ t\geq s>a_{0}\}\subset\mathcal{B}(X)$ an evolution family. We say that $\mathcal{T}$ admits an $h$-dichotomy on an interval $J\subset(a_{0},\infty)$ if there is a family $\{P(t):\ t\in J\}$ of bounded projections on $X$ and constants $D,\lambda>0$ such that:
  1. (1)
    for $t,s\in J$ with $t\geq s$,
    $P(t)T(t,s)=T(t,s)P(s),$ (2.6)
    and
    $T(t,s)\rvert_{\operatorname{Ker}P(s)}\colon\operatorname{Ker}P(s)\to\operatorname{Ker}P(t)\quad\text{is invertible};$ (2.7)
  2. (2)
    for $t,s\in J$ with $t\geq s$,
    $\|T(t,s)P(s)\|\leq D\left(\frac{h(t)}{h(s)}\right)^{-\lambda};$ (2.8)
  3. (3)
    for $t,s\in J$ with $t\leq s$,
    $\|T(t,s)(\text{\rm Id}-P(s))\|\leq D\left(\frac{h(s)}{h(t)}\right)^{-\lambda},$ (2.9)
    where
    $T(t,s):=\left(T(s,t)\rvert_{\operatorname{Ker}P(t)}\right)^{-1}\colon\operatorname{Ker}P(s)\to\operatorname{Ker}P(t).$
Remark 4.
In the particular case where $h(t)=e^{t}$, the notion of a $h$-dichotomy coincides with the notion of an exponential dichotomy as introduced by Henry (see [8, Definition 7.6.1]).
3. Time-rescaling
Throughout this section, we take a growth rate $h\colon(a_{0},\infty)\to(0,\infty)$ and an evolution family $\mathcal{T}=\{T(t,s):\ t\geq s>a_{0}\}\subset\mathcal{B}(X)$. Set
$T_{h}(t,s):=T(h^{-1}(e^{t}),h^{-1}(e^{s})),\quad t,s\in\mathbb{R},\ t\geq s.$
Remark 5.
Observe that $e^{s}>0$ for any $s\in\mathbb{R}$. Consequently, for $s\in\mathbb{R}$, $e^{s}$ belongs to the domain of $h^{-1}\colon(0,\infty)\to(a_{0},\infty)$ and $h^{-1}(e^{s})>a_{0}$. Therefore, since $h^{-1}$ is increasing we have $h^{-1}(e^{t})\geq h^{-1}(e^{s})>a_{0}$ for any $t,s\in\mathbb{R}$ with $t\geq s$. We conclude that $T_{h}(t,s)$ is well-defined.
We then have the following simple observation.
Proposition 1.
$\mathcal{T}^{h}:=\{T_{h}(t,s):\ t\geq s>-\infty\}$ is an evolution family.
Proof.
For $t\in\mathbb{R}$, we have
$T_{h}(t,t)=T(h^{-1}(e^{t}),h^{-1}(e^{t}))=\text{\rm Id}.$
Moreover, for $t\geq s\geq r$ we have
$\begin{split}T_{h}(t,r)=T(h^{-1}(e^{t}),h^{-1}(e^{r}))&=T(h^{-1}(e^{t}),h^{-1}(e^{s}))T(h^{-1}(e^{s}),h^{-1}(e^{r}))\\ &=T_{h}(t,s)T_{h}(s,r).\end{split}$
Here, we used $h^{-1}(e^{t})\geq h^{-1}(e^{s})\geq h^{-1}(e^{r})$ (as $h^{-1}$ is increasing).
Finally, for all $s\in\mathbb{R}$ and $v\in X$, the map $t\mapsto T_{h}(t,s)v$ is continuous on $s,\infty)$ as it is a composition of continuous maps $h^{-1}(e^{s}),\infty)\ni t\mapsto T(t,h^{-1}(e^{s}))v$ and $t\mapsto h^{-1}(e^{t})$. ∎
Remark 6.
We notice that $\mathcal{T}^{h}$ is invertible provided that $\mathcal{T}$ is invertible. Indeed, for $t,s\in\mathbb{R}$ with $t\geq s$, the inverse of $T_{h}(t,s)$ is $T(h^{-1}(e^{s}),h^{-1}(e^{t}))$ (see Remark 1).
Remark 7.
[…]
The main motivation for introducing $\mathcal{T}^{h}$ is the following result.
Proposition 2.
Let $J=[a_{0}^{*},\infty)$ for $a_{0}^{*}>a_{0}$. The following holds:
  1. (a)
    $\mathcal{T}$ exhibits $h$-bounded growth on $J$ if and only if $\mathcal{T}^{h}$ exhibits a bounded growth on $\ln h(a_{0}^{*}),\infty)$;
  2. (b)
    $\mathcal{T}$ admits a $h$-dichotomy on $J$ if and only if $\mathcal{T}^{h}$ admits an exponential dichotomy on $\ln h(a_{0}^{*}),\infty)$.
Proof.
$(a)$ Suppose that $\mathcal{T}$ exhibits $h$-bounded growth on $J$ and let $K,\mu>0$ be such that (2.1) holds. Then
$\|T_{h}(t,s)\|=\|T(h^{-1}(e^{t}),h^{-1}(e^{s}))\|\leq K\left(\frac{h(h^{-1}(e^{t}))}{h(h^{-1}(e^{s}))}\right)^{\mu}=Ke^{\mu(t-s)}$
for $t\geq s\geq\ln h(a_{0}^{*})$, which implies that $h^{-1}(e^{t})\geq h^{-1}(e^{s})\geq a_{0}^{*}$. Hence, $\mathcal{T}^{h}$ exhibits a bounded growth on $\ln h(a_{0}^{*}),\infty)$. The converse implication can be obtained in an analogous manner, noting that
$T(t,s)=T_{h}(\ln h(t),\ln h(s)),\quad t\geq s>a_{0}.$ (3.1)
$(b)$ Suppose that $\mathcal{T}$ admits a $h$-dichotomy on $J$ and let $P(t)$, $t\in J$ and $D,\lambda>0$ be as in Definition 5. Set
$\tilde{P}(t):=P(h^{-1}(e^{t})),\quad t\geq\ln h(a_{0}^{*}).$
We first observe that
$\begin{split}\tilde{P}(t)T_{h}(t,s)&=P(h^{-1}(e^{t}))T(h^{-1}(e^{t}),h^{-1}(e^{s}))\\ &=T(h^{-1}(e^{t}),h^{-1}(e^{s}))P(h^{-1}(e^{s}))\\ &=T_{h}(t,s)\tilde{P}(s),\end{split}$
for $t\geq s\geq\ln h(a_{0}^{*})$. In addition,
$T_{h}(t,s)\rvert_{\operatorname{Ker}\tilde{P}(s)}=T(h^{-1}(e^{t}),h^{-1}(e^{s}))\rvert_{\operatorname{Ker}P(h^{-1}(e^{s}))}\colon\operatorname{Ker}\tilde{P}(s)\to\operatorname{Ker}\tilde{P}(t)$
is invertible.
Secondly, it follows from (2.8) that
$\begin{split}\|T_{h}(t,s)\tilde{P}(s)\|&=\|T(h^{-1}(e^{t}),h^{-1}(e^{s}))P(h^{-1}(e^{s}))\|\\ &\leq D\left(\frac{h(h^{-1}(e^{t}))}{h(h^{-1}(e^{s}))}\right)^{-\lambda}\\ &=De^{-\lambda(t-s)},\end{split}$
for $t\geq s\geq\ln h(a_{0}^{*})$. Similarly, (2.9) gives
$\|T_{h}(t,s)(\text{\rm Id}-\tilde{P}(s))\|\leq De^{-\lambda(s-t)},$
for $\ln h(a_{0}^{*})\leq t\leq s$. We conclude that $\mathcal{T}^{h}$ admits an exponential dichotomy on $\ln h(a_{0}^{*}),\infty)$. The converse implication can be obtained similarly by relying on (3.1). ∎
Remark 8.
Similarly to the proof of Proposition 2(a), one can show that for an invertible evolution family $\mathcal{T}$ the following holds: $\mathcal{T}$ exhibits $h$-bounded decay on $J$ if and only if $\mathcal{T}^{h}$ exhibits bounded decay on $\ln h(a_{0}^{*}),\infty)$.
4. Noncritical uniformity and expansiveness
We introduce the concept of $h$-expansivity for invertible evolution families.
Definition 6.
Let $\mathcal{T}=\{T(t,s):\ t,s>a_{0}\}\subset\mathcal{B}(X)$ be an invertible evolution family and $h\colon(a_{0},\infty)\to(0,\infty)$ be a growth rate. We say that $\mathcal{T}$ is $h$-expansive on an interval $J\subset(a_{0},\infty)$ if there exist $L,\beta>0$ such that
$\|v\|\leq L\left(\left(\frac{h(t)}{h(a)}\right)^{-\beta}\|T(a,t)v\|+\left(\frac{h(b)}{h(t)}\right)^{-\beta}\|T(b,t)v\|\right),$ (4.1)
for $v\in X$ and $a\leq t\leq b$ with $a,b]\subset J$.
Remark 9.
In the case $h(t)=e^{t}$, we will say that $\mathcal{T}$ is exponentially expansive. For evolution families arising from nonautonomous ordinary differential equations, this definition coincides with [15, Definition 5].
Proposition 3.
Let $\mathcal{T}=\{T(t,s):\ t\geq s>a_{0}\}\subset\mathcal{B}(X)$ be an invertible evolution family and $h\colon(a_{0},\infty)\to(0,\infty)$ be a growth rate. Then $\mathcal{T}$ is $h$-expansive on $a_{0}^{*},\infty)$, $a_{0}^{*}>a_{0}$ if and only if $\mathcal{T}^{h}$ is exponentially expansive on $\ln h(a_{0}^{*}),\infty)$.
Proof.
Suppose that $\mathcal{T}$ is $h$-expansive on $a_{0}^{*},\infty)$ and take an arbitrary $c,d]\subset[\ln h(a_{0}^{*}),\infty)$. Let $a_{0}^{*}\leq a\leq b$ be such that $\ln h(a)=c$ and $\ln h(b)=d$. It follows from (4.1) that for any $t\in[c,d]$ and $v\in X$ we have
$\begin{split}\|v\|&\leq L\bigg(\left(\frac{h(h^{-1}(e^{t}))}{h(h^{-1}(e^{c}))}\right)^{-\beta}\|T(h^{-1}(e^{c}),h^{-1}(e^{t}))v\|\\ &\phantom{\leq}+\left(\frac{h(h^{-1}(e^{d}))}{h(h^{-1}(e^{t}))}\right)^{-\beta}\|T(h^{-1}(e^{d}),h^{-1}(e^{t}))v\|\bigg)\end{split},$
as $h^{-1}(e^{t})\in[a,b]=[h^{-1}(e^{c}),h^{-1}(e^{d})]\subset[a_{0}^{*},\infty)$. Hence,
$\|v\|\leq L(e^{-\beta(t-c)}\|T_{h}(c,t)v\|+e^{-\beta(d-t)}\|T_{h}(d,t)v\|).$
We conclude that $\mathcal{T}^{h}$ is exponentially expansive on $\ln h(a_{0}^{*}),\infty)$. The converse implication can be established in an analogous manner on the basis of (3.1).
We now introduce the concept of uniform $h$-noncriticality for evolution families.
Definition 7.
Let $\mathcal{T}=\{T(t,s):\ t\geq s>a_{0}\}\subset\mathcal{B}(X)$ be an invertible evolution family and $h\colon(a_{0},\infty)\to(0,\infty)$ be a growth rate. We say that $\mathcal{T}$ is uniformly $h$-noncritical on $a_{0}^{*},\infty)$ for $a_{0}^{*}>a_{0}$ if there exist $\theta\in(0,1)$ and $C>0$ such that
$\|v\|\leq\theta\sup\{\|T(u,t)v\|:|\ln h(u)-\ln h(t)|\leq C\},$ (4.2)
for all $v\in X$ and $t$ such that $h(t)\geq e^{C}h(a_{0}^{*})$.
Remark 10.
In the case $h(t)=e^{t}$, we will say that $\mathcal{T}$ is uniformly noncritical. For evolution families arising from nonautonomous ordinary differential equations, this definition coincides with [15, Definition 4]. This notion was introduced for nonlinear systems by Krasovski [9] and later adapted to linear systems by Massera and Schäffer [11].
The following result follows easily from the previous definition.
Proposition 4.
Let $\mathcal{T}=\{T(t,s):\ t\geq s>a_{0}\}\subset\mathcal{B}(X)$ be an invertible evolution family and $h\colon(a_{0},\infty)\to(0,\infty)$ be a growth rate. Then $\mathcal{T}$ is uniformly $h$-noncritical on $a_{0}^{*},\infty)$, $a_{0}^{*}>a_{0}$ if and only if $\mathcal{T}^{h}$ is uniformly noncritical on $\ln h(a_{0}^{*}),\infty)$.
Proof.
Assume that $\mathcal{T}$ is uniformly $h$-noncritical and let $\theta\in(0,1)$ and $C>0$ be as in the previous definition. Note that
$\begin{split}&\sup\{\|T_{h}(u,t)v\|:\ |u-t|\leq C\}\\ &=\sup\{\|T(h^{-1}(e^{u}),h^{-1}(e^{t}))v\|:\ |\ln h(h^{-1}(e^{u}))-\ln h(h^{-1}(e^{t}))|\leq C\}.\end{split}$
Consequently, for $t\geq C+\ln h(a_{0}^{*})$ (so that $h(h^{-1}(e^{t}))\geq e^{C}h(a_{0}^{*}))$ and $v\in X$ we have
$\begin{split}\|v\|&\leq\theta\sup\{\|T(s,h^{-1}(e^{t}))v\|:\ |\ln h(s)-\ln h(h^{-1}(e^{t}))|\leq C\}\\ &=\theta\sup\{\|T(h^{-1}(e^{u}),h^{-1}(e^{t}))\|:\ |\ln h(h^{-1}(e^{u}))-\ln h(h^{-1}(e^{t}))|\leq C\}\\ &=\theta\sup\{\|T_{h}(u,t)v\|:\ |u-t|\leq C\}.\end{split}$
We conclude that $\mathcal{T}^{h}$ is uniformly noncritical on $\ln h(a_{0}^{*}),\infty)$. The converse can be established similarly. ∎
5. Main results
5.1. Characterization of exponential dichotomies
The following is the first main result of our paper. Its proof is inspired by the proof of [2, Proposition 1, p.14].
Theorem 1.
Let $\mathcal{T}=\{T(t,s):\ t\geq s>a_{0}\}\subset\mathcal{B}(X)$ be an invertible evolution family and $a_{0}^{*}>a_{0}$ such that the following holds:
  • $\mathcal{T}$ exhibits bounded growth and decay on $a_{0}^{*},\infty)$;
  • there exists a finite-dimensional subspace $Z\subset X$ such that
    $X=\mathcal{S}\oplus Z,$ (5.1)
    where
    $\mathcal{S}:=\left\{v\in X:\ \sup_{t\geq a_{0}^{*}}\|T(t,a_{0}^{*})v\|<+\infty\right\}.$ (5.2)
Then the following statements are equivalent:
  1. (a)
    $\mathcal{T}$ admits an exponential dichotomy on $a_{0}^{*},\infty)$;
  2. (b)
    $\mathcal{T}$ is exponentially expansive on $a_{0}^{*},\infty)$;
  3. (c)
    $\mathcal{T}$ is uniformly noncritical on $a_{0}^{*},\infty)$.
Proof.
$(a)\implies(b)$ Let $P(t)$, $t\geq a_{0}^{*}$ and $D,\lambda>0$ be as in Definition 5 (for $h(t)=e^{t}$). Then, for every $v\in X$ and $a,b]\subset[a_{0}^{*},\infty)$ we have
$v=P(t)v+(\text{\rm Id}-P(t))v=T(t,a)P(a)T(a,t)v+T(t,b)(\text{\rm Id}-P(b))T(b,t)v,$
for $t\in[a,b]$. Consequently, by (2.8) and (2.9) we have
$\begin{split}\|v\|&\leq\|T(t,a)P(a)T(a,t)v\|+\|T(t,b)(\text{\rm Id}-P(b))T(b,t)v\|\\ &\leq De^{-\lambda(t-a)}\|T(a,t)v\|+De^{-\lambda(b-t)}\|T(b,t)v\|,\end{split}$
for $t\in[a,b]$. Hence, (4.1) holds with $\beta=\lambda$, $L=D$ and $h(t)=e^{t}$. We conclude that $\mathcal{T}$ is exponentially expansive on $a_{0}^{*},\infty)$.
$(b)\implies(c)$ Suppose that $\mathcal{T}$ is exponentially expansive on $a_{0}^{*},\infty)$ and let $L,\beta>0$ be such that (4.1) holds with $h(t)=e^{t}$. Choose $C>0$ sufficiently large so that $\theta:=2Le^{-\beta C}<1$. For any $v\in X$ and $t$ such that $t\geq a_{0}^{*}+C$, by (4.1) we have
$\begin{split}\|v\|&\leq L(e^{-\beta C}\|T(t-C,t)v\|+e^{-\beta C}\|T(t+C,t)v\|)\\ &\leq\theta\sup\{\|T(u,t)v\|:\ |u-t|\leq C\},\end{split}$
as $t\in[t-C,t+C]\subset[a_{0}^{*},\infty)$. We conclude that $\mathcal{T}$ is uniformly noncritical on $a_{0}^{*},\infty)$.
$(c)\implies(a)$ Suppose that $\mathcal{T}$ is uniformly noncritical on $a_{0}^{*},\infty)$, and let $C>0$ and $\theta\in(0,1)$ be such that
$\|v\|\leq\theta\sup\{\|T(u,t)v\|:\ |u-t|\leq C\},\quad\text{for $v\in X$ and $t\geq a_{0}^{*}+C$.}$ (5.3)
Moreover, let $K,\mu>0$ be such that (2.1) and (2.2) hold with $h(t)=e^{t}$.
Let
$\mathcal{S}(s):=\left\{v\in X:\ \sup_{t\geq s}\|T(t,s)v\|<+\infty\right\},\quad s\geq a_{0}^{*}.$
Since $T(s,a_{0}^{*})\mathcal{S}=\mathcal{S}(s)$, by (5.1) we have
$X=\mathcal{S}(s)\oplus Z(s),\quad\text{for $s\geq a_{0}^{*}$, where $Z(s):=T(s,a_{0}^{*})Z$.}$ (5.4)
Here $T(s,a_{0}^{*})Z$ denotes the image of the subspace $Z$ under the action of $T(s,a_{0}^{*})$. Fix $s\geq a_{0}^{*}$ and $v\in\mathcal{S}(s)\setminus\{0\}$. Then
$0<\varrho:=\sup_{t\geq s}\|T(t,s)v\|<+\infty.$
For $t\geq s+C$, it follows from (5.3) that
$\begin{split}\|T(t,s)v\|&\leq\theta\sup\{\|T(u,t)T(t,s)v\|:\ |u-t|\leq C\}\\ &\leq\theta\sup\{\|T(u,s)v\|:\ u\geq s\}=\theta\varrho,\end{split}$
yielding
$\sup_{t\geq s+C}\|T(t,s)v\|\leq\theta\varrho<\varrho,$
since $\theta\in(0,1)$. Therefore,
$\varrho=\sup_{t\in[s,s+C]}\|T(t,s)v\|,$
and, consequently, $\varrho\leq Ke^{\mu C}\|v\|$ (where we used (2.1)). We conclude that
$\|T(t,s)v\|\leq D\|v\|\quad\text{for $t\geq s$ and $v\in\mathcal{S}(s)$,}$ (5.5)
where $D\geq 1$ is independent of $s$ and $v$. As in the proof of [2, Proposition 1, p.15], (5.5) (together with (5.3)) implies that
$\|T(t,s)v\|\leq Be^{-\alpha(t-s)}\|v\|\quad\text{for $t\geq s\geq a_{0}^{*}$ and $v\in\mathcal{S}(s)$,}$ (5.6)
where $B:=\theta^{-1}D$ and $\alpha:=-C^{-1}\ln\theta$. Indeed, take $v\in X$, $t\geq s\geq a_{0}^{*}$ and let $n\in\mathbb{N}_{0}$ be such that $s+nC\leq t<s+(n+1)C$. By (5.3),
$\begin{split}\|T(t,s)v\|&\leq\theta\sup\{\|T(u,t)T(t,s)v\|:\ |u-t|\leq C\}\\ &=\theta\sup\{\|T(u,s)v\|:\ |u-t|\leq C\}.\end{split}$ (5.7)
For $u\in[t-C,t+C]$, we have (using (5.3) again) that
$\|T(u,s)v\|\leq\theta\sup\{\|T(r,s)v\|:\ |r-u|\leq C\}$ (5.8)
Noting that $r\in[t-2C,t+2C]$ for each $r$ with $|r-u|\leq C$, we obtain from (5.7) and (5.8) that
$\|T(t,s)v\|\leq\theta^{2}\sup\{\|T(u,s)v\|:\ |u-t|\leq 2C\}.$
Iterating,
$\|T(t,s)v\|\leq\theta^{n}\sup\{\|T(u,s)v\|:\ |u-t|\leq nC\}\leq D\theta^{n}\|v\|,$
where in the last inequality, we used (5.5). This easily implies (5.6).
For $v\in Z$ with $\|v\|=1$, the map $t\mapsto\|T(t,a_{0}^{*})v\|$ is unbounded (as otherwise we would have $v\in\mathcal{S}\cap Z=\{0\}$) and, consequently, there is a least value $t_{1}=t_{1}(v)>a_{0}^{*}$ such that $\|T(t_{1},a_{0}^{*})v\|=\theta^{-1}D$. We claim that $v\mapsto t_{1}(v)$ is bounded. Otherwise, there is a sequence $(v_{n})_{n}\subset Z$ with $\|v_{n}\|=1$ such that $t_{1}(v_{n})\to\infty$. By the compactness of the unit sphere in $Z$ (here we use that $Z$ is finite-dimensional), we can assume without loss of generality that $v_{n}\to v$, where $v\in Z$ and $\|v\|=1$. Since $T(t,a_{0}^{*})\in\mathcal{B}(X)$, we have
$T(t,a_{0}^{*})v_{n}\to T(t,a_{0}^{*})v\quad\text{for $t\geq a_{0}^{*}$.}$
As
$\|T(t,a_{0}^{*})v_{n}\|<\theta^{-1}D\quad\text{for $n\in\mathbb{N}$ and $a_{0}^{*}\leq t<t_{1}(v_{n})$,}$
we have
$\|T(t,a_{0}^{*})v\|=\lim_{n\to\infty}\|T(t,a_{0}^{*})v_{n}\|\leq\theta^{-1}D,\quad t\geq a_{0}^{*}.$
This implies that the map $t\mapsto\|T(t,a_{0}^{*})v\|$ is bounded, which yields a contradiction. This proves our claim. Let $t_{1}:=\sup_{v\in Z,\|v\|=1}t_{1}(v)\in(a_{0}^{*},\infty)$.
As in [2, p.15] we have
$\|T(t,a_{0}^{*})v\|\leq Be^{-\alpha(s-t)}\|T(s,a_{0}^{*})v\|,\quad\text{for $v\in Z$ and $t_{1}\leq t\leq s$.}$
Hence,
$\|T(t,s)v\|\leq Be^{-\alpha(s-t)}\|v\|,\quad\text{for $v\in Z(s)$ and $t_{1}\leq t\leq s$.}$ (5.9)
On the other hand, for $a_{0}^{*}\leq t\leq s\leq t_{1}$ we have from (2.2) that
$\|T(t,s)v\|\leq Ke^{\mu(s-t)}\|v\|\leq Ke^{(\mu+\alpha)(t_{1}-a_{0}^{*})}e^{-\alpha(s-t)}\|v\|,$ (5.10)
for $v\in Z(s)$ as
$e^{\mu(s-t)}=e^{(\mu+\alpha)(s-t)}e^{-\alpha(s-t)}\leq e^{(\mu+\alpha)(t_{1}-a_{0}^{*})}e^{-\alpha(s-t)}.$
Finally, for $a_{0}^{*}\leq t\leq t_{1}\leq s$ using (2.2) and (5.9) we have
$\begin{split}\|T(t,s)v\|=\|T(t,t_{1})T(t_{1},s)v\|&\leq Ke^{\mu(t_{1}-t)}\|T(t_{1},s)v\|\\ &\leq KBe^{\mu(t_{1}-t)}e^{-\alpha(s-t_{1})}\|v\|\\ &=KBe^{(\mu+\alpha)(t_{1}-t)}e^{-\alpha(s-t)}\|v\|\\ &\leq KBe^{(\mu+\alpha)(t_{1}-a_{0}^{*})}e^{-\alpha(s-t)}\|v\|,\end{split}$ (5.11)
for $v\in Z(s)$. From (5.9), (5.10) and (5.11) we conclude that there exists $\tilde{B}>0$ such that
$\|T(t,s)v\|\leq\tilde{B}e^{-\alpha(s-t)}\|v\|,\quad\text{for $v\in Z(s)$ and $a_{0}^{*}\leq t\leq s$.}$ (5.12)
It follows from (2.1), (5.6) and (5.11) (together with [12, Lemma 4.2]) that $\mathcal{T}$ admits an exponential dichotomy on $a_{0}^{*},\infty)$ with respect to projections $P(t)\colon X\to\mathcal{S}(t)$ associated to (5.4). ∎
Remark 11.
  1. (1)
    When $X$ is finite-dimensional, the assumption that there is a finite-dimensional subspace $Z\subset X$ satisfying (5.1) is automatically satisfied.
  2. (2)
    When the evolution family $T(t,s)$ is associated with an ordinary differential equation
    $x^{\prime}=A(t)x,$
    where $A\colon(a_{0},\infty)\to\mathcal{B}(X)$ is continuous, we can remove the assumption that $T(t,s)$ exhibits a bounded decay (see the argument in [2, p.13]).
The following example shows that the assumption that $\mathcal{S}$ has a finite-dimensional complement cannot be omitted in the statement of Theorem 1 (even when $X$ is a Hilbert space).
Example 1.
Let $H=L^{2}(\mathbb{R})$ be the Hilbert space consisting of all square-integrable functions $f\colon\mathbb{R}\to\mathbb{C}$ with respect to the Lebesgue measure on $\mathbb{R}$. For $t\in\mathbb{R}$, let $S(t)\colon H\to H$ be given by
$(S(t)f)(x):=f(x-t),\quad x\in\mathbb{R},\ f\in H.$
Then $S(t)$ is a linear isometry on $H$ for each $t\in\mathbb{R}$. Note that
$S(0)=\text{\rm Id}\quad\text{and}\quad S(t+s)=S(t)S(s)\quad\text{for $t,s\in\mathbb{R}$.}$
Moreover, set $Y:=H\oplus H$ (which is also a Hilbert space). For $t\in\mathbb{R}$, we define $T(t)\colon Y\to Y$ by
$T(t)(f,g)=(e^{-t}S(t)f,e^{t}S(t)g),\quad(f,g)\in Y.$
Clearly, $T(t)$ is a bounded linear operator on $Y$. Let $\mathcal{T}=\{T(t,s):\ t,s\in\mathbb{R}\}$ be an invertible evolution family on $Y$ defined by
$T(t,s):=T(t-s),\quad t,s\in\mathbb{R}.$
We claim that $\mathcal{T}$ admits an exponential dichotomy. To this end, we consider a projection $P\colon Y\to Y$ given by
$P(f,g)=(f,0),\quad(f,g)\in Y,$
and set $P(t):=P$ for $t\in\mathbb{R}$. Clearly, (2.6) holds. Moreover,
$\begin{split}\|T(t,s)P(s)(f,g)\|=\|T(t-s)(f,0)\|&=e^{-(t-s)}\|S(t-s)f\|\\ &=e^{-(t-s)}\|f\|\\ &\leq e^{-(t-s)}\|(f,g)\|,\end{split}$
for $t\geq s$ and $(f,g)\in Y$. Similarly,
$\|T(t,s)(\text{\rm Id}-P(s))(f,g)\|\leq e^{-(s-t)}\|(f,g)\|,$
for $t\leq s$ and $(f,g)\in Y$. We conclude that $\mathcal{T}$ admits an exponential dichotomy.
Next, we consider the operator $M\colon\mathcal{D}(M)\to L^{2}(\mathbb{R})$ defined by
$(Mf)(x)=e^{2x}f(x)\quad x\in\mathbb{R},$
defined on the domain $\mathcal{D}(M)$ consisting of all $f\in L^{2}(\mathbb{R})$ such that $Mf\in L^{2}(\mathbb{R})$. Notice that $\mathcal{D}(M)$ is dense in $L^{2}(\mathbb{R})$ as it contains all continuous functions with bounded support. We observe that $M$ is a closed operator. Indeed, if $(f_{n})_{n}$ is a sequence in $\mathcal{D}(M)$ such that $f_{n}\to f$ and $Mf_{n}\to g$ in $L^{2}(\mathbb{R})$, then we can find a subsequence $(f_{n_{k}})_{k}$ of $(f_{n})_{n}$ such that $f_{n_{k}}\to f$ and $Mf_{n_{k}}\to g$ almost everywhere. This easily implies that $e^{2x}f(x)=g(x)$ for a.e. $x\in\mathbb{R}$. Consequently, $f\in\mathcal{D}(M)$ and $Mf=g$.
We take $X$ to the graph of $M$:
$X=\{(f,Mf):\ f\in\mathcal{D}(M)\}.$
Since $M$ is closed, we have that $X$ is a closed subspace of $Y$. Take $t\in\mathbb{R}$ and $(f,Mf)\in X$ and define $h:=e^{-t}S(t)f$. Note that
$(Mh)(x)=e^{2x}h(x)=e^{2x}e^{-t}(S(t)f)(x)=e^{2x}e^{-t}f(x-t),$
for $x\in\mathbb{R}$. Since $f\in\mathcal{D}(M)$, one can easily see that $Mh\in L^{2}(\mathbb{R})$ and $h\in\mathcal{D}(M)$. Moreover,
$e^{t}(S(t)Mf)(x)=e^{t}e^{2(x-t)}f(x-t)=e^{-t}e^{2x}f(x-t)=Mh(x),$
for $x\in\mathbb{R}$. Hence,
$T(t)(f,Mf)=(e^{-t}S(t)f,e^{t}S(t)Mf)=(h,Mh).$
This implies that $X$ is $T(t,s)$-invariant for arbitrary $t,s\in\mathbb{R}$. Moreover,
$\begin{split}\|T(t)(f,Mf)\|^{2}&=\|(h,Mh)\|^{2}\\ &=\|h\|^{2}+\|Mh\|^{2}\\ &=e^{-2t}\|S(t)f\|^{2}+e^{2t}\|S(t)Mf\|^{2}\\ &=e^{-2t}\|f\|^{2}+e^{2t}\|Mf\|^{2},\end{split}$ (5.13)
for arbitrary $t\in\mathbb{R}$. Note that for $(f,Mf)\in X\setminus\{0\}$ we have that $f\neq 0$ and $Mf\neq 0$. Consequently, (5.13) implies that
$\lim_{t\to\infty}\|T(t)(f,Mf)\|=\infty\quad\text{and}\quad\lim_{t\to-\infty}\|T(t)(f,Mf)\|=\infty,$ (5.14)
for each nonzero $(f,Mf)\in X$. This implies that an invertible evolution family $\bar{\mathcal{T}}=\{\bar{T}(t,s):\ t,s\in\mathbb{R}\}$ on $X$ defined by
$\bar{T}(t,s):=T(t,s)\rvert_{X}\quad t,s\in\mathbb{R}$
does not admit an exponential dichotomy on an interval $0,\infty)$. Indeed, assume that it does admit an exponential dichotomy on $0,\infty)$ with respect to projections $P(t)$, $t\geq 0$. As
$\begin{split}\operatorname{Im}P(t)&=\left\{(f,Mf)\in X:\ \sup_{r\geq t}\|\bar{T}(r,t)(f,Mf)\|<+\infty\right\}\\ &=\left\{(f,Mf)\in X:\ \sup_{r\geq t}\|T(r-t)(f,Mf)\|<+\infty\right\}\\ &=\left\{(f,Mf)\in X:\sup_{r\geq 0}\|T(r)(f,Mf)\|=0\right\},\end{split}$
we conclude from the first equality in (5.14) that $P(t)=0$ for each $t\geq 0$. Consequently, we have that there exist $D,\lambda>0$ such that
$\|\bar{T}(t,s)(f,Mf)\|\geq\frac{1}{D}e^{\lambda(t-s)}\|(f,Mf)\|,$
for each $t\geq s\geq 0$ and $(f,Mf)\in X$. Hence,
$\|T(-r)(f,Mf)\|\leq De^{-\lambda r}\|(f,Mf)\|$
for each $r\geq 0$ and $(f,Mf)\in X$, which contradicts the second equality in (5.14).
Since $\mathcal{T}$ admits an exponential dichotomy, it follows from the proof of Theorem 1 that $\mathcal{T}$ is uniformly noncritical on $0,\infty)$. Thus, there are $C>0$ and $\theta\in(0,1)$ such that
$\|(f,g)\|\leq\theta\sup\{\|T(u,t)(f,g)\|:\ |u-t|\leq C\},\ \text{for $(f,g)\in Y$ and $t\geq C$.}$
In particular,
$\|(f,Mf)\|\leq\theta\sup\{\|T(u,t)(f,Mf)\|:\ |u-t|\leq C\},$
for $(f,Mf)\in X$ and $t\geq C$, yielding
$\|(f,Mf)\|\leq\theta\sup\{\|\bar{T}(u,t)(f,Mf)\|:\ |u-t|\leq C\},$
for $(f,Mf)\in X$ and $t\geq C$. Therefore, $\bar{\mathcal{T}}$ is uniformly noncritical on $0,\infty)$. We conclude that $\bar{\mathcal{T}}$ is uniformly noncritical on $0,\infty)$ although it does not admit an exponential dichotomy on $0,\infty)$.
We finally observe that it follows from (5.14) that $\mathcal{S}$ in (5.2) (with $a_{0}^{*}=0$) corresponding to $\bar{\mathcal{T}}$ consist only of a zero vector, which means that (5.1) holds with $Z=X$. Therefore, we conclude that in the statement of Theorem 1 we cannot relax the assumption that $Z$ is finite-dimensional by requiring only that $Z$ is closed.
5.2. Characterization of $h$-dichotomies
Now we obtain the version of Theorem 1 for $h$-dichotomies. Its proof is a simple consequence of Theorem 1 and the other results of this paper.
Theorem 2.
Let $\mathcal{T}=\{T(t,s):\ t\geq s>a_{0}\}\subset\mathcal{B}(X)$ be an invertible evolution family, $h\colon(a_{0},\infty)\to(0,\infty)$ a growth rate and $a_{0}^{*}>a_{0}$ such that the following holds:
  • $\mathcal{T}$ exhibits $h$-bounded growth and decay on $a_{0}^{*},\infty)$;
  • there exists a finite-dimensional subspace $Z\subset X$ such that
    $X=\mathcal{S}\oplus Z,$
    where
    $\mathcal{S}:=\left\{v\in X:\ \sup_{t\geq a_{0}^{*}}\|T(t,a_{0}^{*})v\|<+\infty\right\}.$
Then the following statements are equivalent:
  1. (a)
    $\mathcal{T}$ admits an $h$-dichotomy on $a_{0}^{*},\infty)$;
  2. (b)
    $\mathcal{T}$ is $h$-expansive on $a_{0}^{*},\infty)$;
  3. (c)
    $\mathcal{T}$ is uniformly $h$-noncritical on $a_{0}^{*},\infty)$.
Proof.
The desired conclusion follows directly from Theorem 1 by noting the following:
  • $\mathcal{T}^{h}$ exhibits bounded growth and decay on $\ln h(a_{0}^{*}),\infty)$ (see Proposition 2(a) and Remark 8);
  • $\mathcal{T}$ admits $h$-dichotomy on $a_{0}^{*},\infty)$ if and only if $\mathcal{T}^{h}$ admits exponential dichotomy on $\ln h(a_{0}^{*}),\infty)$ (see Proposition 2(b));
  • $\mathcal{T}$ is $h$-expansive on $a_{0}^{*},\infty)$ if and only if $\mathcal{T}^{h}$ is exponentially expansive on $\ln h(a_{0}^{*}),\infty)$ (see Proposition 3);
  • $\mathcal{T}$ is uniformly $h$-noncritical on $a_{0}^{*},\infty)$ if and only if $\mathcal{T}^{h}$ is uniformly noncritical on $\ln h(a_{0}^{*}),\infty)$ (see Proposition 4);
  • $\mathcal{S}=\left\{v\in X:\ \sup_{t\geq\ln h(a_{0}^{*})}\|T_{h}(t,\ln h(a_{0}^{*}))v\|<+\infty\right\}.$
Acknowledgements
It is a pleasure to thank all three referees for their careful reading of my paper and for their useful comments. I am also deeply grateful to Prof. M. Pituk for suggesting Example 1.
Funding This paper has been funded by European Union – NextGenerationEU-Statistical properties of random dynamical systems and other contributions to mathematical analysis and probability theory.
Data availability This manuscript has no associated data.
Declarations
Conflict of interest The author declares that he has no known competing financial interests or personal relationships that could have appeared to influence the work in this paper.